OpenStudy (anonymous):

16^-n =8^2n simplify

OpenStudy (p0sitr0n):

add a "log" at each side and solve

OpenStudy (p0sitr0n):

\[(2^4)^(-n) = (2^3)^(2n)\] \[2^(-4n) = 2^(6n)\]

OpenStudy (anonymous):

ok so what would the answer be

OpenStudy (p0sitr0n):

16^-n = 64^n 1/16^n = 64^n n = 0, i suppose for you

OpenStudy (anonymous):

thank you so much!! i got it!!

OpenStudy (p0sitr0n):

youre welcome